1. 100 kmols of moist hydrogen gas containing 4% water by mole is burnt completely in a

furnace with 32% excess dry air.

a) Complete analysis

b) Orsat analysis

2. A furnace is charged with 100 kmols of fuel gas, the composition of which is: 50% CO,

30% H2, 5% CH4, 5% C2H6 and 10% O2. Dry air is supplied 50% in excess for combustion.

Assuming complete combustion.

a) Complete analysis

b) Orsat analysis


1
Expert's answer
2021-02-24T03:46:56-0500

1. 2H2 + O2 = 2H2O

100 x 0.96 = 96 kmols H2.

96/20 x 0.68= 70,5882353 kmols O2.

4 kmols H2O.

2. 2CO + O2 = 2CO2

2H2 + O2 = 2H2O

CH4 + 2O2 = CO2 2+ 2H2O

C2H6 + 2.5O2 = CO2 + 3H2O

50/2= 25 O2. 50 CO2

30 /2= 15 O2. 30 H2O.

5x2 = 10 O2. 5 CO2 and 10 H2O.

5x2 = 10 O2. 5 CO2 and 10 H2O.

25+15+10+10=60 kmols O2

60/0.5 = 120. 120 + 6 = 126 O2